PRELIMINARY MODELING OF CH_{3}D FROM 4000 TO 4550 cm^{−1}

A. V. NIKITIN, Laboratory of Theoretical Spectroscopy, Institute of Atmospheric Optics, Russian Academy of Sciences, 634055 Tomsk, Russian Federation; \underline{LINDA R. BROWN}, K. SUNG, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; M. REY, Vl. G. TYUTEREV, Groupe de Spectromtrie Molculaire et Atmosphrique, UMR CNRS 6089, Universit de Reims, U.F.R. Sciences, B.P. 1039, 51687 Reims Cedex 2, France; M. A. H. SMITH, Science Directorate, NASA Langley Research Center, Hampton, VA 23681, USA; A. W. MANTZ, Dept. of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320, USA.

New line positions, intensities and quantum assignments were obtained for ^{12}CH_{3}D between 4000 and 4550 cm^{−1} using high resolution spectra recorded with two Fourier transform spectrometers and enriched gas samples (98%-Deuterium) at room and cold (80 K) temperatures. The effective Hamiltonian and the effective Dipole moment expansions were expressed in terms of irreducible tensor operators adapted to symmetric top molecules. Over 4000 positions and 1300 line intensities in this region were reproduced with RMS values of 0.007 cm^{−1} and 14%, respectively. With these new results included, the simultaneous modeling of the four lowest polyads of ^{12}CH_{3}D from 0 to 4550 cm^{−1} [Ground State, Triad (1000 1600 cm^{−1}), Nonad (2000 to 3300 cm^{−1}), and Enneadecad (3400 to 4600 cm^{−1})] permitted over 19000 observed positions to be fitted within 0.003 cm^{−1}. ^{a}

aPart of the research described in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, Connecticut College, and NASA Langley Research Center under contracts and cooperative agreements with the National Aeronautics and Space Administration. Support from CH4@Titan contract (ANR France) and of SAMIA GDRI project for the modeling is acknowledged.